TR = 120 Q – Q2

TC = 12 Q2 +30 Q + 10

Determine the quantity level at which the firm maximizes its total profit


1
Expert's answer
2021-04-01T06:57:37-0400

By the definition of the marginal revenue:


"MR=\dfrac{d}{dQ}(TR)=\dfrac{d}{dQ}(120Q-Q^2)=120-2Q."

By the definition the marginal cost:


"MC=\dfrac{d}{dQ}(TC)=\dfrac{d}{dQ}(12Q^2+30Q+10)=24Q+30."

The firm maximizes its total profit when MR=MC:


"120-2Q=24Q+30,""26Q=90,""Q=3.46"